Toss a fair coin thrice. Let x denote the number of tails. Find Var(x)

1
Expert's answer
2022-01-20T15:23:28-0500

Let us now write out all the possible outcomes from tossing 3 coins in a row:

{HHH, HHT, HTH, THH, TTH, THT, HTT, TTT}

Hence n=8

A probability distribution is basically all the possible values of the random variable and their corresponding probabilities. So for our case, this would be the probability distribution:




Variance="\sum x^2*p(x)-\mu^2"

where "\mu=\sum x*p(x)"

="(0*\frac{1}{8})+(1*\frac{3}{8})+(2*\frac{3}{8})+(3*\frac{1}{8})"

"=\frac{3}{2}"

Now we will find the sum :

"\sum x^2*p(x)=(0^2*\frac{1}{8})+(1^2*\frac{3}{8})+(2^2*\frac{3}{8})+(3^2*\frac{1}{8})"

"=3"

Hence variance ="\sum x^2*p(x)-\mu^2"

"=3-(\frac{3}{2})^2"

"=\frac{3}{4}"